OpenStudy (anonymous):

The sum of two numbers is 71. The diffrence is 13, what are the numbers

OpenStudy (anonymous):

Don't give me answer just give me he equation

myininaya (myininaya):

x+y=71 x-y=13

OpenStudy (anonymous):

now eliminate y

OpenStudy (anonymous):

The sum of two numbers is 71 x + y = 71 The difference is 13 x - y = 13

OpenStudy (anonymous):

X = 71 x = 13 and then I...

myininaya (myininaya):

?

OpenStudy (anonymous):

add it 2x = 71+13

OpenStudy (anonymous):

84 dived by two equals 42

OpenStudy (anonymous):

42 and 29

myininaya (myininaya):

yep!

OpenStudy (anonymous):

yea :D

myininaya (myininaya):

gj

OpenStudy (anonymous):

thks